OpenStudy (anonymous):

Can someone help me please to solve this? \[\sum_{k=1}^{n} f _{k} = f _{n+2} = 1\] In this problem \[\f _{k}\] denotes the nth Fibonacci number

OpenStudy (anonymous):

hi can u help me please?

OpenStudy (anonymous):

hi mathmate

OpenStudy (anonymous):

were u able to finish that problem?

OpenStudy (mathmate):

Not sure if you have a typo or not. Have you done proof by induction?

OpenStudy (anonymous):

I don't know how to do that

OpenStudy (mathmate):

I was working on that problem, and tried to absorb the right hand-side into the constants, but I don't get the right answer, so still working on it.

OpenStudy (anonymous):

i made a typo with the \[f _{k}\]...there shouldn't be a backslash before it

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

the instruction is saying to prove \[\sum_{k=1}^{n} f _{k} = f _{n+2} = 1\]

OpenStudy (mathmate):

I see two equal signs, is that correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oh wait

OpenStudy (anonymous):

i think the teacher made a typo

OpenStudy (anonymous):

\[\sum_{k=1}^{n} f _{k} = f _{n+2} - 1\]

OpenStudy (mathmate):

To prove: fi=1,1,2,3,5,8,13,21,... This can be proved by induction. Base case: n=1 \[\sum_{1}^{1} f _{i} = 1\] \[f _{3}-1 = 2-1 =1\] Proposition works for n=1.

OpenStudy (mathmate):

Induction step: Assuming proposition is true for n=k, then \[\sum_{1}^{k} = f _{k+2} - 1\] Then for k+1, add fk+1 to each side to get \[\sum_{1}^{k}i + f _{k+1} = f _{k+2} + f _{k+1} -1\] \[\sum_{1}^{k+1}i = f _{k+3} -1\] QED

OpenStudy (anonymous):

QED?

OpenStudy (mathmate):

"quod erat demonstrandum" latin for what we would like to demonstrate.

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

thank you

OpenStudy (mathmate):

You're welcome. Will get back to the recurrence relation.

OpenStudy (anonymous):

ok no problem :)

OpenStudy (anonymous):

\[\sum_{k=1}^{n} f _{2k-1} = f _{2n}\]

OpenStudy (mathmate):

It's the same prniciple as before: Base case: n=1 \[\sum_{i=1}^{}f _{i} = f _{1} = 1\] \[f _{2*1} = f _{2} =1\] therefore proposition valid for n=1. Assume proposition valid for n=k, then \[\sum_{i=1}^{k}f _{2i-1} = f _{2k}\] add the next term to the left-hand side: \[\sum_{i=1}^{k}f _{2i-1} + f _{2k+1} = \sum_{i=1}^{k+1}f _{2i-1} \] add the same term to the right hand side: \[ f _{2k} + f _{2k+1} = f _{2k+2} \] by definition of the fibonacci sequnce. Therefore the proposition is valid also for n=k+1, therefore the proposition is valid for all integers. QED